Sol Summary – February 11th


Summary Title: Wish List and Microbiology Sample Catalogue

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol summary: The crew had a lovely Canadian/Indian Cultural night, learning about two cultures and getting our names written in Hindi.

Today we woke up and wrote our wish list for the activities we would love to do with our remaining time on Mars. Number one is movie night, games night and salsa night as we keep on postponing it due to the high number of projects we would like to achieve. We still have 4 days and only one cultural night left.

The first EVA in the morning left just 15 min late as one of the wishes from the commander was to take the EVA checklists more seriously. Still the EVA Support team forgot to put a camera in the airlock, which caused the EVA team to do the depressurization twice.

Once on the road, the EVA team managed to install the antenna on the desired location and achieved all their objectives. The rest of the crew was working on finishing the robot and managed to see it turning on for the first time. Unfortunately, the crew could not manage to connect to the robot and has to troubleshoot further tomorrow. For lunch, the team had tortilla bowls with fresh mung bean sprouts that the crew grew over the last couple of days. Mung beans are a powerhouse of vitamins such as A,B,C and E as well as minerals calcium, magnesium, potassium and phosphorous. It also contains zinc, chromium, iron and they taste very fresh and like peas, which is great for Martians in confined space.

Also, today the inherited Kambucha batch of crew 220 should be ready to be drunk by the crew. A first try will be done tomorrow.

During the afternoon, the EVA team got delayed one hour as the crew only had four EVA radios and two of them had been used before and were not recharged in time. In addition, the crew made a mistake in the EVA window and so when the team was finally ready to go, the team was informed that the EVA window is already over. Therefore, the EVA was canceled.

The free time was used to take four crew members’ profile pictures, work on the sample catalogue of the microbiology samples and identify the bacteria based on their appearance such as, form, elevation, surface, opacity and color. The samples from February 3rd were then discarded to avoid any bad bacteria growth. As we are working on fulfilling our wish list, we will have a games night tonight, after a lovely dinner with fresh salad and cucumber from the Greenhab harvested by J. Todd!

Ad Astra!

Anomalies in work: None

Update mouse in hab: Mouse trab reinstalled no new mouse caught

Weather: sunny no winds

Crew Physical Status: Crew status with signs of flu feel better, also crew with knee issues is feeling better.

Reports to be filed:

Sol Summary

Photo Report

Operations Report

GreenHab Report

Science Report

Astronomy Report

EVA Request

EVA Report 10

Journalist Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site